OpenStudy (theconomist):

For my IA, we need to find an experimental value of gravity. I have found the acceleration of a cart on a ramp (0.69m/s^2) - just confused how I find g? (Angle of ramp is 3.3 degrees). Thanks! :-)

OpenStudy (anonymous):



OpenStudy (anonymous):

\[ g = \frac{F_g}{m} \]And \[ a = \frac{F_g\sin(\theta)}{m} \]This means: \[ g\sin(\theta) = a \implies g = a\csc(\theta) \]

OpenStudy (comrad):

You're in IB?

OpenStudy (theconomist):

@Comrad yes